A farmer uses a tractor to pull a 170 kg bale of h
Answer #1


Power output is P = F*v

F is the net force = m*g*sin(15) - fk

v = 5 km/hr = 5*5/18 = 1.38 m/sec

fk is the kinetic friction = mu_k*m*g*cos(15) = (0.4*170*9.8*cos(15)) = 643.7 N

F = (170*9.8*sin(15))-643.7) = -212.5 N


So the tractor has to apply a minimum force of 212.5 N

So power output is P = F*v = 212.5*1.38 =293.25 W